Use the following info for 19-21:

                  i. DTL Jan 1, 2014, $60,000

                  ii. T/I for 2014, $20,000

iii. cumulative temporary difference at Dec 31, 2014 gives rise to future tax amounts $230,000

iv. T/R = 40%

19. J/E for 2014?

                  Dr. DTL

                  Dr. ITP

                  Cr. DTL

                  Cr. ITE

20. ITE on 2014 I/S?

                  8,000

                  32,000

                  40,000

                  108,000

21. Ending DTL balance?

                  8,000

                  32,000

                  92,000

                  108,000
Answer #1

Solution 19:

Required DTL balance at the end of 2014 = $230,000*40% = $92,000

Existing balance of DTL = $60,000

DTL to be created during 2014 = $92,000 - $60,000 = $32,000

J/E for 2014 include credit to DTL

Hence 3rd option is correct.

Solution 20:

Income tax payable for 2014 = $20,000*40% = $8,000

ITE on I/S = $8,000 + $32,000 = $40,000

Hence 3rd option is correct.

Solution 21:

Ending DTL balance = $230,000*40% = $92,000

Hence 3rd option is correct.